OpenStudy (anonymous):

Solve the following system of equations. x + 3y - z = 2 x - 2y + 3z = 7 x + 2y - 5z = -21

OpenStudy (texaschic101):

does it have to be solved a certain way ? substitution, elimination, matrix, or can you solve it any way you choose ?

OpenStudy (anonymous):

Like the answer choices are (2, 3, 5) (-2, 3, 5) (2, -3, 5) (2, 3, -5)

OpenStudy (anonymous):

do you have a graphing calculator?..

OpenStudy (texaschic101):

x + 3y - z = 2 --->(-1) x - 2y + 3z = 7 --------------- -x - 3y + z = -2 (result of multiplying by -1) x - 2y + 3z = 7 ---------------add -5y + 4z = 5 x - 2y + 3z = 7 --->(-1) x + 2y - 5z = -21 --------------- -x + 2y - 3z = -7 (result of multiplying by -1) x + 2y - 5z = -21 ---------------add 4y - 8z = - 28 -5y + 4z = 5 --->(2) 4y - 8z = -28 -------------- -10y + 8z = 10 (result of multiplying by 2) 4y - 8z = -28 -------------add -6y = -18 y = -18/-6 y = 3 4y - 8z = -28 4(3) - 8z = -28 12 - 8z = -28 -8z = -28 - 12 -8z = - 40 z = -40/-8 z = 5 x + 3y - z = 2 x + 3(3) - 5 = 2 x + 9 - 5 = 2 x + 4 = 2 x = 2 - 4 x = -2 solution is : (-2,3,5)